# AT Loss: Advanced Torrential Loss Function for Precipitation Forecasting

### v1

@@ Abstract
Table 2: Forecast performance comparisons.
@@
2020-minute lead time
@@
$40$-minute lead time
@@
$60$-minute lead time
@@
To enable a more intuitive comparison, Fig. 1 illustrates a case of $1$-hour forecasts at 01:50:00 UTC on June $26$, $2024$. The highlighted areas represent regions with precipitation intensity greater than or equal to $2$ mm/h. In this case, the AT loss produced the best forecasts beginning from the $30$-minute lead time. In particular, it most accurately simulated the eastward-moving precipitation echoes. As a supplement, an additional case is presented in Fig. 3 of the End Matter.
Consistency Experiments under Outliers.—The performance consistency of the AT loss under outliers was verified through comparative experiments with baselines. To use only the intended outliers in the experiments, the interquartile range of the original data was first calculated, and Tukey’s method was applied to identify the existing outliers within the range [31, 32]. The identified outliers were replaced with the mean precipitation intensity of neighboring grid cells. Then, this refined data was used to train a convolutional neural network (CNN) for $1$-step forecasting in two tracks, clean and dirty, for each loss function, with architectural details provided in the End Matter. The clean track used the refined data as is for training. In the dirty track, intentional noise was injected as the refined data was fed into the CNN. Specifically, $10$ to $30$ percent of the refined data was randomly corrupted with salt-and-pepper noise and random-valued impulse noise [33, 34]. After training, forecasts from both tracks were compared in terms of MAE and peak signal-to-noise ratio (PSNR) [35, 36]. The lower the MAE and the higher the PSNR between the forecasts from the clean and dirty tracks, the more consistent the performance of the loss function is, regardless of contamination.
Table 3: Performance differences under outliers.
@@
Since the primary objective of AT loss is not to determine the exact value of each grid cell but to identify whether it is greater than or equal to the threshold, it is most effective when combined with other loss functions that guide the model toward accurate values. Therefore, when applying AT loss in operational models, it is important to seek a balance with these complementary loss functions. In other words, empirical strategies aimed at achieving optimal forecasts, including adjusting the weights across loss functions, employing staged training, and fine-tuning, are essential.
@@ Acknowledgements.
Acknowledgments.—The authors thank Yura Kim and Seongmook Kim for insightful discussions. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)], and the Development of Integrated Application Technology for Korea Weather Radar project funded by the Weather Radar Center, Korea Meteorological Administration [KMA2021-03122, Development of Radar Based Severe Weather Nowcasting Technology]. Code Availability.—The official implementation of the AT loss is publicly available at https://github.com/kaetsraeb/at-loss.

### v2

# Advanced Torrential Loss Function for Precipitation Forecasting

@@ Abstract
For the convenience of scholars from diverse fields, a glossary is also provided in the End Matter.
@@
Table 2: Forecast performance comparisons. The row and column headers indicate loss functions and metrics, respectively. All abbreviations are defined in the text.
@@
22-step forecast ($20$-minute)
@@
44-step forecast ($40$-minute)
@@
66-step forecast ($60$-minute)
@@
To enable a more intuitive comparison, Fig. 1 illustrates a case of $1$-hour forecasts at 01:50:00 UTC on June $26$, $2024$. The highlighted areas represent regions with precipitation intensity greater than or equal to $2$ mm/h. In this case, the AT loss produced the best forecasts beginning from the $30$-minute lead time. In particular, it most accurately simulated the eastward-moving precipitation echoes. As a supplement, Figs. 3 and 4 in the End Matter present extended results of Fig. 1 and an additional case, respectively.
Consistency Experiments under Outliers.—The performance consistency of the AT loss under outliers was verified through comparative experiments with baselines. To use only the intended outliers in the experiments, the interquartile range of the original data was first calculated, and Tukey’s method was applied to identify the existing outliers within the range [31]. The identified outliers were replaced with the mean precipitation intensity of neighboring grid cells. Then, this refined data was used to train a convolutional neural network (CNN) for $1$-step forecasting in two tracks, clean and dirty, for each loss function, with architectural details provided in the End Matter. The clean track used the refined data as is for training. In the dirty track, intentional noise was injected as the refined data was fed into the CNN. Specifically, $10$ to $30$ percent of the refined data was randomly corrupted with salt-and-pepper noise and random-valued impulse noise [32, 33]. After training, forecasts from both tracks were compared in terms of MAE and peak signal-to-noise ratio (PSNR) [34, 35]. The lower the MAE and the higher the PSNR between the forecasts from the clean and dirty tracks, the more consistent the performance of the loss function is, regardless of contamination.
Table 3: Performance differences under outliers. The row and column headers indicate loss functions and metrics, respectively. All abbreviations are defined in the text.
@@
As an extension of these ablation studies, the End Matter discusses in detail the considerations for applying the AT loss to worldwide operational models used in professional institutions.
@@ Acknowledgements.
Acknowledgments.—The authors thank Yura Kim and Seongmook Kim for insightful discussions. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00459387), and the Development of Integrated Application Technology for Korea Weather Radar project funded by the Weather Radar Center, Korea Meteorological Administration [KMA2021-03122, Development of Radar Based Severe Weather Nowcasting Technology]. Code Availability.—The official implementation of the AT loss is publicly available at https://github.com/kaetsraeb/at-loss. Data Availability.—The data that support the findings of this work are openly available [36].